OpenStudy (anonymous):

sqrt(x+iy) =+- (a+ib) , then sqrt(-x-iy) is equal to a. +-(ai -b) b. +_(a-ib)

jimthompson5910 (jim_thompson5910):

hint: \[\Large \sqrt{-x-iy}=\sqrt{-1(x+iy)}=\sqrt{-1}\sqrt{x+iy}=i\sqrt{x+iy}\] So \[\Large \sqrt{-x-iy}=i\sqrt{x+iy}\]

OpenStudy (anonymous):

pls help...i can't see the full text..

jimthompson5910 (jim_thompson5910):

which part is cut off? (sry for the long delay)

OpenStudy (anonymous):

its okie...thanks for helpin.. :)

OpenStudy (anonymous):

i got the answer..